# Reinforcement Learning based Autonomous Multi-Rotor Landing on Moving Platforms

### v2

@@ 5.4.2 Evaluation of the RL Approach
Furthermore, during the real flight experiments no significant jittering in the agent’s actions could be observed. This substantiates the approach of calculating values for the maximum pitch angle and agent frequency presented in Section 3.5.
@@ 5.5 Ablation Study
In the context of our approach and in addition to the presented method where the landing task was learned by means of a sequential curriculum leveraging transfer learning between curriculum steps, there are two other ways of how learning the task could be achieved. First, the sequential curriculum is conducted without transfer learning. This means that whenever a new curriculum step is added to the learning sequence the Q-table associated with the latest curriculum step is initialized with zeros. We denote this ablation case Sequential curriculum without transfer learning. And second, we abandon curriculum learning completely. Instead, we create the full sequence of learning steps right in the beginning with all associated Q-tables initialized with zeros. We then conduct one training on the full sequence without adding an additional step after each round of training. This constitutes a scenario in which several separate Q-learning tasks are trained simultaneously. For this reason, we call this ablation case Simultaneous training of sequence. For the training and evaluation, we chose the scenario RPM 1.6. All learning related parameters are the same as presented in Section 4.5. We compare these additional two methods with the agent selected for further evaluation in Section 5.2 and name this scenario Sequential curriculum with transfer learning. Table 13 summarizes the results.
@@
Simultaneous training of sequence $81$ $86$ $77$ $74$ $72$ $72$ $147$ $2567$
@@
It is clear that ablation case Sequential curriculum without transfer learning performs worst across all evaluation scenarios in terms of both training duration as well as number of episodes required. Our rationale here is the following. In each new curriculum step the task has to be learned from scratch. However, since in all curriculum steps added to the sequence after the initial step the exploration rate is set to $\varepsilon=0$, it is even more difficult to learn optimal behaviour. Furthermore, in the early phase after adding a new curriculum step its behaviour is bad due to few updates of the state-action-pairs, therefore acting as a source of disturbance to the previous curriculum steps. For this reason, in order to reach a situation in which the Q-values of the latest curriculum step as well as the Q-values of the previous curriculum steps have sufficiently adapted ($96\%$ success rate over the last $100$ episodes), a significant amount of time is required. The discrepancy between the success rate necessary to end the training and the success rate achieved during evaluation can be explained as follows. As outlined in Section 3.4, an episode is terminated with success if the agent spends one second in the latest curriculum step before reaching its goal state. However, during evaluation the multi-rotor vehicle is required to constantly stay above the moving platform, not only for one second. This leaves room for erratic behaviour after this one second period which can be observed for this ablation case. Even though centering the multi-rotor vehicle above the platform is learned in principle, maintaining that centered state for a longer time period is not possible which leads to low success rates. Conceivably, the situation could be improved by choosing a longer time period for maintaining the goal state before the success criterion is met.

### v3

@@ 5.4.2 Evaluation of the RL Approach
Furthermore, during the real flight experiments no significant jittering in the agent’s actions could be observed. This substantiates the approach of calculating values for the maximum pitch angle and agent frequency presented in Sect. 3.5.
@@ 5.5 Ablation Study
In the context of our approach and in addition to the presented method where the landing task was learned by means of a sequential curriculum leveraging transfer learning between curriculum steps, there are two other ways of how learning the task could be achieved. First, the sequential curriculum is conducted without transfer learning. This means that whenever a new curriculum step is added to the learning sequence the Q-table associated with the latest curriculum step is initialized with zeros. We denote this ablation case Sequential curriculum without transfer learning. And second, we abandon curriculum learning completely. Instead, we create the full sequence of learning steps right in the beginning with all associated Q-tables initialized with zeros. We then conduct one training on the full sequence without adding an additional step after each round of training. This constitutes a scenario in which several separate Q-learning tasks are trained simultaneously. For this reason, we call this ablation case Simultaneous training of sequence. For the training and evaluation, we chose the scenario RPM 1.6. All learning related parameters are the same as presented in Sect. 4.5. We compare these additional two methods with the agent selected for further evaluation in Sect. 5.2 and name this scenario Sequential curriculum with transfer learning. Table 13 summarizes the results.
@@
Simultaneous training of sequence $81$ $86$ $77$ $74$ $72$ $72$ $147$ $1538$
@@
It is clear that ablation case Sequential curriculum without transfer learning performs worst across all evaluation scenarios in terms of both training duration as well as number of episodes required. Our rationale here is the following. In each new curriculum step the task has to be learned from scratch. However, since in all curriculum steps added to the sequence after the initial step the exploration rate is set to $\varepsilon=0$, it is even more difficult to learn optimal behaviour. Furthermore, in the early phase after adding a new curriculum step its behaviour is bad due to few updates of the state-action-pairs, therefore acting as a source of disturbance to the previous curriculum steps. For this reason, in order to reach a situation in which the Q-values of the latest curriculum step as well as the Q-values of the previous curriculum steps have sufficiently adapted ($96\%$ success rate over the last $100$ episodes), a significant amount of time is required. The discrepancy between the success rate necessary to end the training and the success rate achieved during evaluation can be explained as follows. As outlined in Sect. 3.4, an episode is terminated with success if the agent spends one second in the latest curriculum step before reaching its goal state. However, during evaluation the multi-rotor vehicle is required to constantly stay above the moving platform, not only for one second. This leaves room for erratic behaviour after this one second period which can be observed for this ablation case. Even though centering the multi-rotor vehicle above the platform is learned in principle, maintaining that centered state for a longer time period is not possible which leads to low success rates. Conceivably, the situation could be improved by choosing a longer time period for maintaining the goal state before the success criterion is met.